solve
log4(4b+14)-log4(b^2-3b-17)=1/2

Answer by edjones(7569)

log4(4b+14)-log4(b^2-3b-17)=1/2
log[4]((4b+14)/(b^2-3b-17))=1/2
4^log[4]((4b+14)/(b^2-3b-17))=4^(1/2)
(4b+14)/(b^2-3b-17)=2
4b+14=2b^2-6b-34
2b^2-10b-48=0
b^2-5b-24=0
(b-8)(b+3)=0
b=8, b=-3
.
Ed